Materials Variances Cinturon Corporation produces high-quality leather belts. The company's plant in Boise uses a standard costing system and has set the following standards for materials and labor: Leather (3 strips @ $4) $12.00 Direct labor (0.75 hr. @ $12) 9.00 Total prime cost $21.00

During the first month of the year, the Boise plant produced 92,000 belts. Actual leather purchased was 290,500 strips at $3.00 per strip. There were no beginning or ending inventories of leather. Actual direct labor was 79,000 hours at $12.50 per hour. Required: 1. Break down the total variance for materials into a price variance and a usage variance using the columnar and formula approaches.

Enter favorable values as negative numbers and unfavorable values as positive numbers.

Price variance

Usage variance

Total variance $
Answer #1

Price variance = (standard price - actual price)*actual quantity purchased

= (4-3)*290500 = $290500 Favorable

Usage variance = (standard quantity - actual quantity)*standard price

= (92000*3 - 290500)*4 = $58000 unfavorable

Total variance = 290500-58000 = $232500 Favorable